### v1

@@ 4.1 Datasets
Figure 3: Clef is evaluated on two real-world domains involving multivariate trajectories: (a) cellular development and (b) patient health. (a) To study cellular development, fibroblast cells derived from mice can be artificially reprogrammed into various other cell states in vitro. A cell’s state is defined by its gene expression. Throughout reprogramming, a cell activates transcription factor (TF) genes at different time points to change its gene expression, thereby influencing its developmental trajectory. In this illustration, a mouse fibroblast is being reprogrammed over the span of 20 days (D0-D20); color and shape represent cell state. On day 8, if the cell activates the Obox6 TF, the cell is on the path toward becoming an induced pluripotent stem cell (iPSC); whereas if it activates the Neurod4 TF, it is on the path toward becoming a neuron or astrocyte. (b) The health of a human patient is often monitored through lab tests (e.g. blood sodium level, white blood cell count). The history of lab results across multiple patient visits (V1-V9) as well as candidate clinical interventions (e.g., medication) can be used to infer the most likely future trajectory of the patient’s health. Illustrations from NIAID NIH BIOART Source (see References).
@@
Table 1: Dataset statistics. We construct three core datasets: WOT (cellular developmental trajectories), eICU (patient lab tests), and MIMIC-IV (patient lab tests). We also construct a paired counterfactual cellular trajectories dataset, WOT-CF. $N$ is the number of sequences (i.e., cellular developmental trajectories, patient lab test trajectories), $V$ is the number of measured variables (i.e., gene expression, lab test), and $L$ is the length of the sequences.
@@
4.2 Setup
Metrics. We use standard metrics (MAE, RMSE, and $R^{2}$) to quantify sequence editing performance.
Baselines. We evaluate Clef against a traditional multivariate time series algorithm, Vector Autoregression (VAR) model Lütkepohl (2005). As Clef can leverage any type of sequence encoder, we benchmark against the state-of-the-art condition-guided counterfactual sequence generation setup with different sequential data encoders: Transformer Waswani et al. (2017); Narasimhan et al. (2024); Jing et al. (2024); Zhang et al. (2023a) and xLSTM Beck et al. (2024). We further evaluate Clef against a state-of-the-art time series foundation model, MOMENT Goswami et al. (2024); specifically, we finetune an adapter for the $1024$-dimensional embeddings generated by the frozen MOMENT-1-large embedding model.
Ablations. To investigate the effectiveness of the learned temporal concepts, we evaluate against an ablated model, SimpleLinear, in which temporal concepts are simply all ones; in other words, temporal concepts are not learned nor meaningful. This ablation is inspired by traditional linear models that excel when $\mathbf{x}_{t_{j}}\simeq\mathbf{x}_{t_{i}}$ Toner & Darlow (2024); Ahlmann-Eltze et al. (2024). We also evaluate different versions of Clef with and without an FFN layer in the concept encoder $E$ (Appendix C).
Implementation details. Models are trained on a single NVIDIA A100 or H100 GPU. All models have comparable number of parameters as their Clef-based counterparts. Refer to Appendix B for details and hyperparameter selection.
5 Results
Figure 4: Benchmarking the performance of Clef, baselines, and ablation models on (a) immediate and (b) delayed sequence editing. The models are trained using a standard cell- or patient-centric random split. Not shown for visualization purposes are the performances of VAR models on eICU and MIMIC-IV datasets: on immediate sequence editing, MAE for eICU and MIMIC-IV are $55982.74$ and $886.05$, respectively; on delayed sequence editing, MAE for eICU and MIMIC-IV are $3.02\times 10^{39}$ and $8.62\times 10^{23}$, respectively.
We evaluate Clef’s performance on controllable sequence editing across multiple datasets and tasks. We aim to answer the following research questions. R1: How well does Clef perform in immediate sequence editing? R2: How well does Clef perform in delayed sequence editing? R3: How does Clef generalize to unseen/new sequences? R4: Can Clef perform zero-shot counterfactual generation? R5: How can Clef be leveraged for real-world counterfactual patient trajectory simulations? We establish that Clef outperforms state-of-the-art baselines in sequence editing, demonstrating both immediate and delayed sequence editing capabilities, strong generalizability, and real-world applicability.
5.1 R1: Immediate sequence editing
Immediate sequence editing involves forecasting the next time step of a sequence under a counterfactual condition. This is useful in settings where interventions take effect instantaneously, such as introducing a genetic perturbation in cellular systems or administering a drug to a patient (Definition 3.1). Example counterfactual scenarios in which immediate sequence editing is applicable are: What if we treat the cells with the candidate drug now? and What if we perform surgery on the patient today?
Clef models consistently outperform baseline models across all datasets (Figure 4a; Appendix Figures 8-9). The SimpleLinear baseline, which assumes minimal temporal changes, performs comparably in some cases, but Clef outperforms it on datasets where short-term dynamics are more complex. On WOT, all Clef models outperform or perform comparably to the time series forecasting model, VAR. This is particularly exciting given recent findings that linear models can achieve competitive or better forecasting performance than neural network models Toner & Darlow (2024); Ahlmann-Eltze et al. (2024). These results highlight Clef’s ability to accurately modify trajectories at the right points while preserving unaffected portions of the sequence, an advantage in counterfactual reasoning.
Regardless of the sequence encoder used with Clef, these models tend to outperform or perform comparably to non-Clef models (Figure 4a). However, the performance of Clef can be affected by the ability of the sequence encoder to capture the temporal dynamics of the input sequences. For instance, models with the MOMENT encoder yield the highest MAE in all three datasets, with and without help from Clef (Figure 4a). Nevertheless, Clef models with the MOMENT encoder reduce the MAE of non-Clef models.
5.2 R2: Delayed sequence editing
Delayed sequence editing requires forecasting a counterfactual trajectory at a future time step while maintaining causal consistency. This task is challenging, as small errors can compound over longer horizons. Example scenarios in which delayed sequence editing is applicable are: What if we treat the cells with the candidate drug in ten days? and What if we perform the surgery on the patient next year?
Clef outperforms or performs competitively against SimpleLinear and VAR on the patient datasets, eICU and MIMIC-IV (Figure 4b; Appendix Figures 8-9). Clef-transformer and Clef-xLSTM achieve lower MAE than SimpleLinear, whereas non-Clef transformer and MOMENT baselines perform comparably or worse. As in immediate sequence editing, models using MOMENT as the sequence encoder (i.e., using temporal concepts with the MOMENT sequence encoder) yield the highest MAE. However, incorporating Clef with MOMENT reduces the MAE to levels comparable to SimpleLinear and VAR.
On WOT, SimpleLinear and VAR outperform neural network models in delayed sequence editing (Figure 4b). This suggests that cellular developmental trajectories exhibit small and possibly noisy changes at each time step, favoring linear models Ahlmann-Eltze et al. (2024); Toner & Darlow (2024). Additionally, given the relatively small number of training trajectories compared to the high-dimensional state space, nonlinear models may overfit to noise more readily than linear models. Nevertheless, Clef significantly reduces the MAE of non-Clef models, demonstrating its effectiveness as a regularizer that mitigates short-term noise while preserving long-term trends.
5.3 R3: Generalization to new patient trajectories
We assess the ability of Clef models to generalize to new patient sequences. To evaluate robustness, we use the SPECTRA approach Ektefaie et al. (2024) to create challenging data splits where the test sets have minimal similarity to the training data (Appendix A.2).
Across both the eICU and MIMIC-IV patient datasets, Clef models exhibit stronger generalization than non-Clef models (Figure 5; Appendix Figures 11-12 and Table 2). For immediate and delayed sequence editing on eICU, Clef-transformer and Clef-xLSTM maintain stable and strong performance even as train/test divergence increases. In contrast, their non-Clef counterparts degrade significantly. Although baseline MOMENT models show relatively stable performance across train/test splits in delayed sequence editing, they generalize poorly compared to Clef-MOMENT models. Despite similar performance between xLSTM and Clef-xLSTM in delayed sequence editing on both patient datasets (Figure 4b), Clef-xLSTM demonstrates superior generalizability (Figure 5b), highlighting the effectiveness of Clef in adapting to unseen data distributions.
Figure 5: Generalizability of Clef on (a) eICU and (b) MIMIC-IV patient datasets in immediate and delayed sequence editing. As the SPECTRA parameter increases, the train/test split similarity decreases (Appendix Figure 10). The area under the spectral performance curve (AUSPC) evaluation is in Appendix Table 2.
5.4 R4: Zero-shot counterfactual generation of cellular trajectories
In addition to evaluating Clef’s generalizability to new patient lab test trajectories, we assess on zero-shot counterfactual generation for cellular trajectories (Figure 6; Appendix Figure 13). Using the Waddington-OT model, we generate sequences that remain consistent until a specified divergence time step, where an alternative condition—such as the activation of a different transcription factor—introduces a shift (Appendix A.1). This process yields $1,273$ pairs of “original” and “counterfactual” trajectories, totaling $2,546$ individual sequences (Table 1). Models are trained on the “original” trajectories and evaluated on the “counterfactual” trajectories in a zero-shot setting.
Clef-based models consistently outperform non-Clef models in both immediate and delayed sequence editing (Appendix Figure 13). To more closely analyze delayed sequence editing performance, we examine the predictions for cellular trajectories of length 23, the most common sequence length in the dataset (Figure 6). Since $t_{i}=10$ is the earliest divergence time step, we provide the first nine time steps $\mathbf{x}_{:,0:9}$, the counterfactual condition, and $t_{j}\in[10,23]$ to the model. Comparing the generated and ground truth counterfactual sequences, we find that Clef significantly outperforms non-Clef models after time step 10, which is when the trajectories begin to diverge (Figure 6).
Figure 6: Zero-shot counterfactual generation of cellular developmental trajectories. Shown are the MAE of predictions at each time step for counterfactual sequences of length 23 (the most common sequence length in the dataset) starting at time step 10 (the earliest divergence time step of a counterfactual trajectory).
5.5 R5: Case studies using real-world patient datasets
We evaluate Clef’s ability to simulate counterfactual patient trajectories through temporal concept intervention. We conduct case studies on two independent cohorts of patients with type 1 diabetes mellitus (T1D), a chronic autoimmune disease in which the immune system attacks insulin-producing cells in the pancreas Quattrin et al. (2023).
Unlike counterfactual generation methods that rely on condition tokens to guide generation Narasimhan et al. (2024); Jing et al. (2024); Zhang et al. (2023a), Clef allows direct edits to the generated outputs to produce counterfactual sequences. This capability is particularly valuable when condition tokens are insufficient, such as when prescribing medication dosage. Instead of relying on predefined conditions, Clef can precisely modify the values of specific lab tests to explore their longitudinal effects.
Setup. For an individual patient, we intervene on the temporal concepts corresponding to specific lab tests to simulate the “reversal” or “worsening” of symptoms, thereby generating “healthier” or “more severe” trajectories, respectively. Formally, given temporal concept $\mathbf{c}$ learned from $\mathbf{x}_{:,t_{0}:t_{i}}$ and an optional condition $s$, we modify $\mathbf{c}^{I}\neq\mathbf{c}$ such that at least one element satisfies $\mathbf{c}_{k}\neq\mathbf{c}^{I}_{k}$.
From the eICU and MIMIC-IV datasets, we construct two independent cohorts of T1D patients and matched healthy individuals (Appendix A.2). The eICU-T1D dataset contains $59$ T1D patients and $579$ matched healthy controls, while MIMIC-IV-T1D includes $25$ T1D patients and $226$ matched healthy controls.
To generate counterfactual sequences, we modify specific values in temporal concept $\mathbf{c}$, such as glucose levels, and allow Clef to simulate future trajectories of length $T=10$. We then compare these counterfactual trajectories (i.e., Clef-generated patients) against observed sequences from matched healthy individuals, other healthy individuals, and other T1D patients. Our hypothesis is that clinically meaningful edits will produce “healthier” (i.e., more similar to healthy patients) or “sicker” (i.e., more similar to other T1D patients) trajectories.
Results. First, we modify Clef’s concepts to reduce glucose levels by half, aligning them closer to normal physiological ranges. The resulting counterfactual patient trajectories exhibit higher $R^{2}$ similarity with both matched and other healthy individuals compared to other T1D patients (Figure 7a). This suggests that Clef effectively generates counterfactual trajectories indicative of a healthier state.
Next, we simulate a worsening condition by doubling glucose levels. The resulting counterfactual trajectories generated by Clef show higher $R^{2}$ similarity with other T1D patients than with healthy individuals (Figure 7a), as would be expected based on clinical evidence.
Beyond direct interventions, we examine indirect changes in Clef-generated patients’ lab values resulting from glucose modifications. In both eICU-T1D and MIMIC-IV-T1D cohorts, lowering glucose also leads to a reduction in white blood cell (WBC) count (Figure 7b; Appendix Figure 14a). This aligns with clinical knowledge, as T1D is an autoimmune disorder where immune activity, including WBC levels, plays a critical role Quattrin et al. (2023). Further, when we intervene on Clef to reduce WBC levels instead of glucose, we observe a concurrent drop in glucose across both cohorts (Appendix Figure 14b,c), reinforcing the interconnected nature of these physiological markers.
Figure 7: Clef-generated counterfactual patients via intervention on temporal concepts. We intervene on Clef to (a) halve (top) or double (bottom) a T1D patient’s glucose levels to infer a “healthier” or “sicker” counterfactual patient, respectively. (b) Observed and Clef patients from the eICU-T1D cohort are compared to quantify the differences between their lab test trajectories (indirect effects) as a result of the intervention to halve T1D patients’ glucose levels.
Finally, we demonstrate that modifying multiple lab tests simultaneously can produce compounding effects. When we intervene on Clef to reduce both glucose and WBC levels, the resulting Clef-generated patients resemble healthy individuals even more closely than other T1D patients (Appendix Figure 14d). This finding suggests that Clef can integrate multiple simultaneous edits, capturing their joint impact on a patient’s future state.
6 Conclusion
In this work, we formalize controllable sequence editing for counterfactual generation on biomedical sequences, and demonstrate that Clef outperforms state-of-the-art counterfactual sequence generation models in immediate and delayed sequence editing. Clef also has stronger generalizability to new sequences, and performs significantly better than state-of-the-art models in zero-shot counterfactual generation. Further, we show that interventions directly on Clef’s temporal concepts can generate counterfactual patients such that their trajectories are shifted toward healthier outcomes. This capability has the potential to help discover clinical interventions that could alleviate a patient’s symptoms. While this work focuses on cellular and patient trajectories, Clef can be readily extended to perform sequence editing in other domains.
Limitations. There are two key limitations of Clef. Firstly, we define temporal concepts such that each element represents a unique measured variable in the sequence (e.g., gene expression, lab test). Instead, it may be beneficial to learn higher-order relationships between the measured variables or across time as abstract hierarchical concepts The et al. (2024); Kacprzyk et al. (2024). Secondly, while Clef is able to generate counterfactual sequences for any condition, including those it may not have seen during training, Clef could potentially improve with additional guidance from a real-world causal model for the system or domain of interest Chatzi et al. (2024). Since defining such a real-world causal graph is a major challenge, one promising future direction could be to enable user interventions, such as those performed in our T1D case studies, to finetune Clef.
Broader Impact
By introducing a flexible and interpretable approach to counterfactual sequence generation, Clef bridges the gap between language model-style conditional generation and structured, time-sensitive sequence editing, with implications for decision support in medical and scientific applications. Like all generative AI models, Clef (and its derivatives) should be used solely for the benefit of society. In this study, we demonstrate that Clef can generate alternative cellular trajectories and simulate the reversal or progression of symptoms to model healthier or sicker patient outcomes. However, this work (and any derivatives) should never be used to induce harmful cellular states (e.g., activating transcription factors to drive a cell toward a pathological state) or negatively impact patient care (e.g., neglecting necessary clinical interventions or recommending harmful treatments). Our goal is to help researchers understand the underlying mechanisms of disease to improve public health. Any misuse of this work poses risks to patient well-being. Therefore, the ability to intervene on Clef’s generated outputs should be leveraged to assess the model’s robustness and correctness for ethical and responsible use.
Acknowledgements
M.L. and M.Z. are supported by the Berkowitz Family Living Laboratory at Harvard Medical School and the Clalit Research Institute. Y.E. is supported by grant T32 HG002295 from the National Human Genome Research Institute and the NDSEG Fellowship. We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, ARPA-H BDF program, awards from Chan Zuckerberg Initiative, Bill & Melinda Gates Foundation INV-079038, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, Harvard Data Science Initiative, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
Appendix A Data & Experimental Setup
In this section, we provide further details about data construction, data preparation, and experimental setup. We share code and instructions in our GitHub repository to reproduce the experiments in this paper: https://github.com/mims-harvard/CLEF.
A.1 Cellular Developmental Trajectories
@@ A.1.1 Simulating trajectories
Cellular reprogramming experiments help elucidate cellular development Schiebinger et al. (2019). In these experiments, cells are manipulated and allowed to progress for a specific period of time before they undergo RNA sequencing (RNA-seq), and we analyze the resulting RNA-seq data to observe their new cellular profiles Schiebinger et al. (2019). RNA-seq is a destructive process for the cell, meaning that the same cell cannot be sequenced at two different time points. Computational models are thus necessary to infer the trajectory of a cell.
@@
Waddington-OT dataset and model. Waddington-OT Schiebinger et al. (2019) is a popular approach to reconstruct the landscape of cellular reprogramming using optimal transport (OT). There are two components in Waddington-OT: (1) a single-cell RNA-seq (scRNA-seq) dataset of mouse cells from a reprogramming experiment, and (2) an OT-based trajectory inference model fitted on the scRNA-seq dataset. The scRNA-seq dataset consists of 251,203 mouse cells profiled from 37 time points (0.5-day intervals) during an 18-day reprogramming experiment starting from mouse embryonic fibroblasts. The trajectory inference model consists of transport matrices $\pi_{t_{k},t_{k+1}}$ with dimensions $N\times M$ that relate all cells $\mathbf{x}^{1}_{t_{k}},...,\mathbf{x}^{n}_{t_{k}}$ profiled at time $t_{k}$ to all cells $\mathbf{x}^{1}_{t_{k+1}},...,\mathbf{x}^{m}_{t_{k+1}}$ profiled at time $t_{k+1}$. An entry at row $i$ and column $j$ of $\pi_{t_{k},t_{k+1}}$ corresponds to the probability that $\mathbf{x}^{j}_{t_{k+1}}$ is a descendant cell of $\mathbf{x}^{i}_{t_{k}}$, as determined using optimal transport Chizat et al. (2017). Every cell in the scRNA-seq dataset is either pre-labeled as one of the 13 provided cell sets (i.e., induced pluripotent stem, stromal, epithelial, mesenchymal-epithelial transition, trophoblast, spongiotrophoblast, trophoblast progenitor, oligodenrocyte progenitor, neuron, radial glial, spiral artery trophoblast giant, astrocyte, other neural) or unlabeled. We cluster the unlabeled cells using Leiden clustering via scanpy Wolf et al. (2018) at a resolution of 1, and define the resulting 27 unlabeled clusters as unique cell sets. As a result, each cell in the dataset belongs one and only one cell set.
@@
Simulating cell state trajectories. We define “cell state” as the transcriptomic profile of a cell. Here, a transcriptomic profile is the log-normalized RNA-seq counts of the top $1,479$ most highly variable genes. To create a simulated trajectory of cell states for an individual cell undergoing reprogramming, we randomly and uniformly sample a cell profiled at time step $t_{0}$ (Day $0.0$) from the Waddington-OT scRNA-seq dataset, and generate via the transport matrix $\pi_{t_{0},t_{1}}$ a probability distribution $\mathbb{P}_{t_{1}}$ over possible descendant cells $\mathbf{x}^{1}_{t_{1}},\dots,\mathbf{x}^{m}_{t_{1}}$ at time step $t_{1}$ (Day $0.5$). We sample a cell from this distribution, and repeat the process until we reach either Day $18.0$ or a terminal state (i.e., neural, stromal, or induced pluripotent stem cell). After generating a trajectory composed of cells from the Waddington-OT scRNA-seq dataset through this process, we retrieve the transcriptomic profile of each cell to compose $\mathbf{x}_{:,t_{0}:t_{T}}$, where $T$ is the length of the trajectory.
@@
Inferring conditions. A condition $s_{t_{i}}$ is defined as the activation of a transcription factor (TF) that leads a cell to transition from state $\mathbf{x}_{t_{i}}$ to descendant state $\mathbf{x}_{t_{i+1}}$. To infer such conditions, we perform differential expression analysis between cells from the same cell set as $\mathbf{x}_{t_{i}}$ (i.e., $\mathbf{x}^{a}\in A$) and cells from the same cell set as $\mathbf{x}_{t_{i+1}}$ (i.e., $\mathbf{x}^{b}\in B$). Using the wot.tmap.diff_exp function (via the Waddington-OT library), we identify the top TF that was significantly upregulated in $\mathbf{x}^{a}\in A$ compared to $\mathbf{x}^{b}\in B$. If no TFs are differentially expressed, then the condition is “None.” We retroactively perform this analysis on all pairs of consecutive cell states in a cell state trajectory $\mathbf{x}_{:,t_{0}:t_{T}}$ to obtain the full trajectory containing both cell states and TF conditions: $\tau=\{\mathbf{x}_{t_{0}},s_{t_{0}},\mathbf{x}_{t_{1}},s_{t_{1}},\cdots,s_{t_{T-1}},\mathbf{x}_{t_{T}}\}$. In other words, $\tau$ represents a simulated trajectory of an individual cell undergoing the reprogramming process. Condition embeddings $\mathbf{z}_{s}\in\mathbb{R}^{5120}$ are obtained from the (frozen) pretrained ESM-2 embedding model Lin et al. (2022).
@@
Generating matched counterfactual trajectories. We additionally create pairs of matched counterfactual trajectories to evaluate a model’s performance in zero-shot counterfactual generation. Each pair consists of an “original” trajectory $\tau_{og}$ and a “counterfactual” trajectory $\tau_{cf}$. First, we generate $\tau_{og}$ using the Waddington-OT model. Then, given a divergence time step $D$, the first $D$ time steps of $\tau_{og}$ are carried over to $\tau_{cf}$ such that the first $D$ cell states and conditions of $\tau_{og}$ and $\tau_{cf}$ are exactly the same. The remaining states and conditions of $\tau_{cf}$ are sampled independently from $\tau_{og}$, resulting in an alternative future trajectory based on an alternative condition at time step $D$.
@@
Implementation note: Because Clef learns time embeddings based on the year, month, date, and hour of a given timestamp, we convert the time steps of each cell into timestamps. We set the starting time $t_{0}$ as timestamp 2000/01/01 00:00:00, and add $10\times t_{i}$ hours to the converted timestamp of $t_{i-1}$.
A.1.2 Experimental setup

### v2

@@ Appendix C Data & Experimental Setup
Overview of novel datasets. To study cellular development, fibroblast cells derived from mice can be artificially reprogrammed into various other cell states in vitro. A cell’s state is defined by its gene expression. Throughout reprogramming, a cell activates transcription factor (TF) genes at different time points to change its gene expression, thereby influencing its developmental trajectory. In Figure 3a, a mouse fibroblast is being reprogrammed over the span of 20 days (D0-D20); color and shape represent cell state. On day 8, if the cell activates the Obox6 TF, the cell is on the path toward becoming an induced pluripotent stem cell (iPSC); whereas if it activates the Neurod4 TF, it is on the path toward becoming a neuron or astrocyte. The health of a human patient is often monitored through lab tests (e.g. blood sodium level, white blood cell count). As shown in Figure 3b, the history of lab results across multiple patient visits (V1-V9) as well as candidate clinical interventions (e.g., medication) can be used to infer the most likely future trajectory of the patient’s health.
@@
Table 1: Dataset statistics for conditional sequence generation benchmarks. We construct three core datasets for benchmarking conditional sequence generation: WOT (cellular developmental trajectories), eICU (patient lab tests), and MIMIC-IV (patient lab tests). We also construct a paired counterfactual cellular trajectories dataset, WOT-CF. $N$ is the number of sequences (i.e., cellular developmental trajectories, patient lab test trajectories), $V$ is the number of measured variables (i.e., gene expression, lab test), and $L$ is the length of the sequences.
@@
C.1 Cellular Developmental Trajectories
@@ C.1.1 Simulating trajectories
Cellular reprogramming experiments help elucidate cellular development [schiebinger2019-ie]. In these wet-lab experiments, cells are manipulated and allowed to progress for a specific period of time before they undergo RNA sequencing (RNA-seq), and we analyze the resulting RNA-seq data to observe their new cellular profiles [schiebinger2019-ie]. RNA-seq is a destructive process for the cell, meaning that the same cell cannot be sequenced at two different time points. Computational models are thus necessary to infer the trajectory of a cell.
@@
Waddington-OT dataset and model. Waddington-OT [schiebinger2019-ie] is a popular approach to reconstruct the landscape of cellular reprogramming using optimal transport (OT). There are two components in Waddington-OT: (1) a single-cell RNA-seq (scRNA-seq) dataset of mouse cells from a reprogramming experiment, and (2) an OT-based trajectory inference model fitted on the scRNA-seq dataset. The scRNA-seq dataset consists of 251,203 mouse cells profiled from 37 time points (0.5-day intervals) during an 18-day reprogramming experiment starting from mouse embryonic fibroblasts. The trajectory inference model consists of transport matrices $\pi_{t_{k},t_{k+1}}$ with dimensions $N\times M$ that relate all cells $\mathbf{x}^{1}_{t_{k}},...,\mathbf{x}^{n}_{t_{k}}$ profiled at time $t_{k}$ to all cells $\mathbf{x}^{1}_{t_{k+1}},...,\mathbf{x}^{m}_{t_{k+1}}$ profiled at time $t_{k+1}$. An entry at row $i$ and column $j$ of $\pi_{t_{k},t_{k+1}}$ corresponds to the probability that $\mathbf{x}^{j}_{t_{k+1}}$ is a descendant cell of $\mathbf{x}^{i}_{t_{k}}$, as determined using optimal transport [chizat2017scalingalgorithm]. Every cell in the scRNA-seq dataset is either pre-labeled as one of the 13 provided cell sets (i.e., induced pluripotent stem, stromal, epithelial, mesenchymal-epithelial transition, trophoblast, spongiotrophoblast, trophoblast progenitor, oligodenrocyte progenitor, neuron, radial glial, spiral artery trophoblast giant, astrocyte, other neural) or unlabeled. We cluster the unlabeled cells using Leiden clustering via scanpy [wolf2018-tb] at a resolution of 1, and define the resulting 27 unlabeled clusters as unique cell sets. As a result, each cell in the dataset belongs one and only one cell set.
@@
Simulating cell state trajectories. We define “cell state" as the transcriptomic profile of a cell. Here, a transcriptomic profile is the log-normalized RNA-seq counts of the top $1,479$ most highly variable genes. To create a simulated trajectory of cell states for an individual cell undergoing reprogramming, we randomly and uniformly sample a cell profiled at time step $t_{0}$ (Day $0.0$) from the Waddington-OT scRNA-seq dataset, and generate via the transport matrix $\pi_{t_{0},t_{1}}$ a probability distribution $\mathbb{P}_{t_{1}}$ over possible descendant cells $\mathbf{x}^{1}_{t_{1}},\dots,\mathbf{x}^{m}_{t_{1}}$ at time step $t_{1}$ (Day $0.5$). We sample a cell from this distribution, and repeat the process until we reach either Day $18.0$ or a terminal state (i.e., neural, stromal, or induced pluripotent stem cell). After generating a trajectory composed of cells from the Waddington-OT scRNA-seq dataset through this process, we retrieve the transcriptomic profile of each cell to compose $\mathbf{x}_{:,t_{0}:t_{T}}$, where $T$ is the length of the trajectory.
@@
Inferring conditions. A condition $s_{t_{i}}$ is defined as the activation of a transcription factor (TF) that leads a cell to transition from state $\mathbf{x}_{t_{i}}$ to descendant state $\mathbf{x}_{t_{i+1}}$. To infer such conditions, we perform differential expression analysis between cells from the same cell set as $\mathbf{x}_{t_{i}}$ (i.e., $\mathbf{x}^{a}\in A$) and cells from the same cell set as $\mathbf{x}_{t_{i+1}}$ (i.e., $\mathbf{x}^{b}\in B$). Using the wot.tmap.diff_exp function (via the Waddington-OT library), we identify the top TF that was significantly upregulated in $\mathbf{x}^{a}\in A$ compared to $\mathbf{x}^{b}\in B$. If no TFs are differentially expressed, then the condition is “None." We retroactively perform this analysis on all pairs of consecutive cell states in a cell state trajectory $\mathbf{x}_{:,t_{0}:t_{T}}$ to obtain the full trajectory containing both cell states and TF conditions: $\tau=\{\mathbf{x}_{t_{0}},s_{t_{0}},\mathbf{x}_{t_{1}},s_{t_{1}},\cdots,s_{t_{T-1}},\mathbf{x}_{t_{T}}\}$. In other words, $\tau$ represents a simulated trajectory of an individual cell undergoing the reprogramming process. Condition embeddings $\mathbf{z}_{s}\in\mathbb{R}^{5120}$ are obtained from the (frozen) pretrained ESM-2 embedding model [lin2022language].
@@
Generating matched counterfactual trajectories. We additionally create pairs of matched counterfactual trajectories to evaluate a model’s performance in zero-shot counterfactual generation. Each pair consists of an “original" trajectory $\tau_{og}$ and a “counterfactual" trajectory $\tau_{cf}$. First, we generate $\tau_{og}$ using the Waddington-OT model. Then, given a divergence time step $D$, the first $D$ time steps of $\tau_{og}$ are carried over to $\tau_{cf}$ such that the first $D$ cell states and conditions of $\tau_{og}$ and $\tau_{cf}$ are exactly the same. The remaining states and conditions of $\tau_{cf}$ are sampled independently from $\tau_{og}$, resulting in an alternative future trajectory based on an alternative condition at time step $D$.
@@
Implementation note: Since Clef learns time embeddings based on the year, month, date, and hour of a given timestamp, we convert the time steps of each cell into timestamps. We set the starting time $t_{0}$ as timestamp 2000/01/01 00:00:00, and add $10\times t_{i}$ hours to the converted timestamp of $t_{i-1}$.
C.1.2 Experimental setup